The gravitational force between two lagre space ships in space may be expressed as Fg = Gr^-2
solve the equation Fg = Gr^-2 for r

Answer by stanbon(75887)

solve the equation Fg = Gr^-2 for r
------
Fg = G/r^2
---
r^2 = G/Fg
---
r = sqrt[G/Fg]
=====================
cheers,
Stan H.